# Breaking the Curse of Dimensionality with Distributed Neural Computation

### v1

@@ 3.1 Preliminaries
Notation. We use the following notation: for any $f,g:\mathbb{R}\to\mathbb{R}$, we write $f\in\mathcal{O}(g)$ if there exist $x_{0}\in\mathbb{R}$ and $M\geq 0$ such that for each $x\geq x_{0}$ we have $|f(x)|\leq Mg(x_{0})$. Similarly, we write $f\in\Omega(g)$ to denote the relation $g\in O(f)$. The ReLU activation function is given for every $x\in\mathbb{R}$ by $\text{ReLU}(x)=\max\{x,0\}$. For each $n\in\mathbb{N}_{+}$ and $C\subseteq\mathbb{R}^{n}$, the indicator function $I_{C}$ of $C$ is defined by: for each $x\in\mathbb{R}^{n}$ set $I_{C}(x)=1$ if $x\in C$ and is 00 otherwise.
@@
Multi-Layer Perceptrons. In this paper, we will consider MLPs with trainable PReLU activation functions.
@@
$\theta$ $\leftrightarrow\big{(}A^{(j)},b^{(j)},\bar{\alpha}^{(j)}\big{)}_{j=0}^{J}$ (3)
$(A^{(j)},b^{(j)},\bar{\alpha}^{(j)})$ $\in\mathbb{R}^{d_{j+1}\times d_{j}}\times\mathbb{R}^{d_{j}}\times\mathbb{R}^{d_{j}}$
With the above identification, we recursively define the representation function of a $[d]$-dimensional network by
@@
$x^{(j+1)}$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\sigma_{\bar{\alpha}^{(j)}}\bullet(A^{(j)}\,x^{(j)}+b^{(j)})$
@@
We now formally define the classes of neural pathways.
Definition 3.2 (Neural Pathways).
Let $J,W,L,n\in\mathbb{N}$ and fix an activation function $\sigma\in C(\mathbb{R})$. The set of neural pathways with at-most $L$ leaves, depth $J$, and width $W$, denoted by $\mathcal{NP}_{J,W,L:n,m}^{\sigma}$, consists of all functions $\hat{f}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ satisfying $\hat{f}=\sum_{i=1}^{l}\,f_{i}\,I_{C_{i}}$ where for $f_{1},\dots,f_{l}\in\mathcal{NN}^{\operatorname{PReLU}}_{J,W}$, and distinct prototypes $p_{1},\dots,p_{L}\in\mathbb{R}^{n}$; whose induced Voronoi cells $\{C_{i}\}_{i=1}^{L}$ are defined recursively by
$C_{i}$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\tilde{C}_{i}\setminus\bigcup_{j<i}\,\tilde{C_{j}}$ (5)
where for $i=1,\dots,l$ the (non-disjoint) cells are
$\tilde{C}_{i}$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\big{\{}x\in[0,1]^{n}:\,\|x-p_{i}\|=\min_{j\in\{1,\dots,l\}}\,\|x-p_{j}\|\big{\}}$
Remark 3.3 (Trees).
The structure in any neural pathway is any tree with root node $\mathbb{R}^{n}$ and leaves given by the pairs $\{(p_{i},f_{i})\}_{i=1}^{l}$ or equivalently $\{(C_{i},f_{i})\}_{i=1}^{l}$. The purpose of any such tree is simply to efficiently route an input $x\in\mathbb{R}^{n}$ to one of the $l$ “leaf networks” $f_{1},\dots,f_{l}$ using $\mathcal{O}(l\log(l))$ queries; to identify which Voronoi cells $\{C_{i}\}_{i=1}^{l}$ the point $x$ is contained in.
@@
shatters it in the classical sense; here $I_{(0,\infty)}(t)=1$ if $t\geq 0$ and equals to 00 otherwise. Denoted by $\operatorname{VC}(\mathcal{F})$, the VC dimension of $\mathcal{F}$ is the cardinality of the largest $k$-point subset shattered by $\mathcal{F}$. If $k$ is unbounded, then we say that $\mathcal{F}$ has an infinite VC dimension (over $\mathcal{X}$). One can show, see Bartlett et al. (2019), that the VC-dimension of any such $\mathcal{F}$ is roughly the same as the pseudo-dimension of Pollard (1990) for a small modification of $\mathcal{F}$.
@@
3.2 Main Theoretical Results
To reduce notation, our main result is formulated for locally-Hölder function on the closed unit ball of $\mathbb{R}^{n}$, defined by $\overline{B}_{n}(0,1)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{x\in\mathbb{R}^{n}:\,\|x\|\leq 1\}$.
Theorem 3.4 (Neural Pathways Avoid The Curse of Dimensionality).
Fix $0<\alpha\leq 1$, dimensions $n,m\in\mathbb{N}$, parameters $0<\alpha\leq 1$ and an $\alpha$-Hölder map $f:\overline{\operatorname{B}}_{n}(0,1)\to\mathbb{R}^{m}$ with optimal Hölder coefficient $L\geq 0$, and $\sigma\in C(\mathbb{R})$ satisfying Definition 3.1.
@@ 3.2 Main Theoretical Results
Fix an “approximation error” $\varepsilon>0$. There is a binary tree $\mathcal{T}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(V,E)$ with leaves =$\mathcal{L}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{(v_{i},\theta_{i})\}_{i=1}^{L}\subseteq\overline{\operatorname{B}}_{n}(0,1)\times\mathbb{R}^{P([d])}$ satisfying
$\max_{x\in\overline{\operatorname{B}}_{n}(0,1)}\,\min_{(v_{i},\theta_{i})\in\mathcal{L}}\,\|x-v_{i}\|\in\mathcal{O}\big{(}\varepsilon^{1/\alpha-1/n}\big{)}.$ (8)
Furthermore, for each $x\in\overline{\operatorname{B}}_{n}(0,1)$, if $v_{i}$ is such that $\|x-v_{i}\|\in\mathcal{O}(\varepsilon^{1/\alpha-1/n})$ then
$\|f(x)-\hat{f}_{\theta_{i}}(x)\|<\varepsilon.$ (9)
The depth and width of each $\hat{f}_{\theta_{i}}$ are recorded in Table 1 and the binary tree $\mathcal{T}$ has $2^{\mathcal{O}(n-\frac{n}{\alpha}\log(\varepsilon))}$  leaves.
Remark 3.5.
A technical version of Theorem 3.4 with $\nu$-ary trees and general uniformly continuous target functions, namely Theorem A.5, can be found in Appendix A.2.
@@
Next, we demonstrate that the neural pathways model can generalize and generate functions that are PAC-learnable, thanks to its finite VC dimension. This property, however, breaks down in MLP models with super-expressive activation functions.
Proposition 3.6 (Neural Pathways Can Generalize).
@@
We complement the main result of Bartlett et al. (2019) by demonstrating that the class of unstable MLPs (Shen et al., 2022a) possesses infinite VC dimension, even when they have finite depth and width. Thus, while they may serve as a gold standard from the perspective of approximation theory, they should not be considered a benchmark gold standard from the viewpoint of learning theory.
@@
Let $\mathcal{F}$ denote the set of MLPs with activation function in Definition A.1, depth at-most $15$, and width at-most $36n(2n+2)$. Then $\operatorname{VC}(\mathcal{F})=\infty$.
@@
The VC dimension bounds for the standard MLP model, MLP with a super-expressive activation function as proposed by Shen et al. (2022a), and the neural pathways model are summarized in Table 2.
@@
Table 2: VC Dimension of the Neural Pathways, ReLU MLP, and MLP model with Super-Expressive Activation function of Shen et al. (2022a). All models have depth $J$, width $W$, and (when applicable) $L$ leaves; where $J,W,L,n\in\mathbb{N}_{+}$.
Model VC Dim Ref. Neural Pathways $\mathcal{O}\big{(}L\log(L)^{2}\,\max\{nL\log(L),JW^{2}\log(JW)\}\big{)}$ Prop 3.6 ReLU MLP $\mathcal{O}\big{(}JW^{2}\log(JW)\big{)}$ Bartlett et al. (2019) Super-Expressive $\infty$ Prop 3.7
4 Related Work and Discussion
With the advent of large foundation models, the use of distributed computing to scale deep learning models beyond what is feasible to process on a single machine has become exceedingly relevant. However, most of the literature focuses on distributed or parallelized GPU training of standard neural network models (Dean et al., 2012; Hegde & Usmani, 2016; Sergeev & Del Balso, 2018; Ben-Nun & Hoefler, 2019). Next, we review relevant literature.
Prototypes. Each region in our learned partition of the input space is associated with a representative point therein called a prototype. Prototypes (also called landmarks) are routinely used in image classification (Mensink et al., 2012), few-shot learning (Snell et al., 2017; Cao et al., 2020), dimensionality reduction (Law et al., 2019), in complex networks (Keller-Ressel & Nargang, 2023), and geometric deep learning (Ghadimi Atigh et al., 2021) to tractably encode massive structures. They are also standard in classical clustering algorithms such as $K$-medoids or $K$-means, wherein the part associated with each medoid (resp. centroid) defines a Voronoi cell or Voronoi region (Voronoi, 1908). In our proposed approach, one key difference is that we optimize prototypes to maximize performance on the loss function defining the learning problem, rather than to divide up the input space based on a clustering objective, which may not optimize the model’s performance. Nevertheless, in principle, many of these prototype-based clustering algorithms can be substituted into our pipeline.
Partitioning. While partitioning is commonly employed in deep learning for various purposes, such as proving universal approximation theorems (Yarotsky, 2017; Lu et al., 2021b; Gühring & Raslan, 2021) or facilitating clustering-based learning (Kratsios & Zamanlooy, 2022; Trask et al., 2022; Ali & Nouy, 2023; Srivastava et al., 2022), existing approaches typically involve loading the entire model into VRAM. Our approach, however, differs by relying on a learned partition of the input space, where each part is associated with a distinct small neural network. Importantly, the complete set of networks forming the neural pathways does not need to be simultaneously loaded into VRAM during training or inference.
Cluster-then-Predict. Several deep learning models employ a cluster-then-predict approach to upscaling, e.g., (Wang et al., 2021). In these approaches, one typically maps an input to a specific prototype and passes that prototype through a relatively small network, effectively reducing dimension by discretizing the input space to consist only of the set of prototypes. This is in stark contrast to the neural pathways framework, which performs no discretization and only hierarchically partitions the input space. Every input is passed through a neural network at the leaf node, corresponding to the nearest prototypes, but the input itself is never replaced by one of the prototypes. Srivastava et al. (2022) also use multiple local networks, but they employ soft clustering for the final prediction, which means they have to load all the local models for a single sample.
Comparison to Standard Distributed Training. One can distribute the complexity of feedforward models by storing each of their layers in offline memory and then loading them sequentially into VRAM during the forward pass. This does avoid loading more than $\mathcal{O}(\operatorname{Width})$ active parameters into VRAM at any given time, where $\operatorname{Width}$ denotes the width of the feedforward model. However, doing so implies that all the model parameters are ultimately loaded during the forward pass. This contrasts with the neural pathways model, which requires $\mathcal{O}\big{(}\log_{2}(K)\,\operatorname{Width}^{2}\,\operatorname{Depth}\big{)}$ to be loaded into memory during a forward pass; where $\operatorname{Width}$ and $\operatorname{Depth}$ are respectively the largest width and depth of the MLP at any leaf of the tree defining a given neural pathway, and $K$ denotes the number of prototypes. However, in the forward pass, one loads $\mathcal{O}(\varepsilon^{-n/2})$ parameters for the best worst-case MLP while only $\mathcal{O}(n\log(1/\varepsilon)/\varepsilon)$ are needed in the case of the neural pathways. The number of parameters here represents the optimal worst-case rates for both models (see Table 1 and Theorem A.5).
Logarithmic number of queries via trees. For many prototypes, as in our main guarantee, the neural pathways only need to evaluate the distance between the given input and a logarithmic number of prototypes—specifically, one for each level in the tree—when using deep binary trees to hierarchically refine the Voronoi cells. Thus, a given machine never processes the exponential number of prototypes, and only $\nu\lceil\log_{\nu}(K)\rceil$ prototypes are ever queried for any given input; when trees are $\nu$-ary (as in Theorem A.5), and where $K$ denotes the number of prototypes. Since we consider that prototypes are queried separately and before loading neural pathways, we do not take them into account when counting the number of learnable parameters. Moreover, the size of our prototypes is negligible in our experiments.
5 Experiments
We evaluate our approach in two standard machine learning tasks: regression and classification. We experimentally show that neural pathways, which distribute predictions over multiple neural networks, are competitive with a single large neural network containing as many model parameters as all the neural pathways combined. This is desirable in cases where the large neural network does not fit into the memory of a single machine. On the contrary, the neural pathways framework can be trained by distributing each neural pathway on separate machines (or equivalently serially on a single machine). Inference can then be performed by loading only a single neural pathway at a time into the GPU. Our source code to reproduce our results is available in the supplementary material.
5.1 Regression
We first consider regression, where the goal is to approximate non-convex synthetic functions often used for performance test problems. In particular, we choose 1-dimensional Hölder functions, as well as Ackley (Ackley, 1987) and Rastrigin (Rastrigin, 1974) functions, whose formulations are detailed in Appendix D.
1D Hölder Functions. We illustrate our primary finding, as encapsulated in Theorem 3.4, by leveraging $1$-dimensional functions characterized by very low regularity. This choice is motivated by the jagged structure inherent in such functions, necessitating an exponentially higher sampling frequency compared to smooth functions for achieving an accurate reconstruction. Indeed, this crucial sampling step forms the foundation of many quantitative universal approximation theorems (Yarotsky, 2017; Shen et al., 2021; Kratsios & Papon, 2022). As elaborated in Appendix C, approximating a well-behaved (Lipschitz) function in $d$ dimensions poses a challenge equivalent to approximating a highly irregular function ($1/d$-Hölder) in a single dimension. A visual representation of a $1/d$-Hölder function is presented in Figure 2, exemplified by the trajectory of a fractional Brownian motion with a Hurst parameter of $\alpha=1/d$. A formal definition is available in Appendix C.
1D Hölder 2D Ackley 3D Ackley 2D Rastrigin 3D Rastrigin
Neural pathways (ours) 0.057 $\pm$ 0.085 0.00015 $\pm$ 0.00006 0.00068 $\pm$ 0.00010 0.0480 $\pm$ 0.0073 1.0062 $\pm$ 0.0446
(a) Ground Truth Ackley
(b) Ground Truth Rastrigin
(c) Predicted Ackley
(d) Predicted Rastrigin
(e) log MSE
(f) log MSE
During training, for a given and fixed set of $K$ prototypes $p\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(p_{1},\dots,p_{K})$, we assign each training sample $x$ to its nearest prototype $p_{k}$ and associated neural network $\hat{f}_{k}$. We learn the prototypes as explained in Section 2.2. For simplicity, we set the number of prototypes to $K=4$; we also set $s=10,000$ if $n=1$, $s=150$ if $n=2$ (i.e., $500150^{2}=22,$ 500 samples), and $s=30$ if $n=3$ (i.e., $27,000$ samples). More details can be found in Appendix D.
Test performance. In Table 3, we present the mean squared error obtained on the test set across 10 random initializations and various splits of the training/test sets. The baseline consists of a single neural network with the same overall architecture as each neural pathway but possesses as many parameters as all the neural pathways combined. In all instances, the neural pathways model demonstrates a significant performance advantage compared to the baseline. Figure 3 illustrates the predictions generated by our neural pathways, showcasing the capability of our approach to achieve a good approximation of the ground truth functions.
5.2 Classification
Training. Our neural pathways model takes as input latent DINOv2 encodings (Oquab et al., 2023) of images from the aforementioned datasets. Each sample $x\in\mathbb{R}^{768}$ corresponds to a DINOv2 embedding (i.e., $n=768$). Additionally, we set the prototypes as centroids obtained through the standard $K$-means clustering on the DINOv2 embedding space. The replacement of our original prototype learning algorithm is sensible in this context, as we operate within a structured latent space optimized through self-supervised learning using large-scale compute and datasets.
Dataset CIFAR-10 CIFAR-100 Food-101
Ours (weighted) 98.40 $\pm$ 0.05 90.01 $\pm$ 0.11 91.86 $\pm$ 0.10
Ours (unweighted) 98.42 $\pm$ 0.04 89.62 $\pm$ 0.25 91.79 $\pm$ 0.16
Baseline 98.45 $\pm$ 0.06 89.85 $\pm$ 0.17 91.45 $\pm$ 1.09
Due to the potential class imbalance in the various Voronoi cells formed by the prototypes, we utilize two variations of the cross-entropy loss for each neural pathway $\hat{f}_{k}$. The first variation, termed unweighted, assigns equal weight to all categories. The weighted variation assigns a weight that is inversely proportional to the distribution of each category in the Voronoi cell defined by the prototype $p_{k}$.
Test performance. We present the test classification accuracy over 10 random initializations of both the baseline and our neural pathways in Table 4. The weighted version performs slightly better on datasets with a large number of categories. Our approach achieves comparable results with the baseline, effectively decomposing the prediction across multiple smaller models.
6 Conclusion
We introduced a novel approach to overcome the curse of dimensionality based on distributed neural computation. Instead of relying on exceptionally regular target functions or super-expressive MLPs, as seen in other related works in the literature, we propose distributing computation across several small specialized networks. These networks can learn to model the target function within Voronoi cells of the input space. Our approach is particularly relevant, given the computational limitations of current AI hardware, handicapped by finite VRAM. This problem has recently been exacerbated by the growing trend to increase model size with large foundation models. In particular, our neural pathways framework can scale up simply by adding more neural networks (or equivalently by refining the Voronoi cells which subdivide the input space), since each independent network can always be loaded into VRAM while the overall model keeps increasing in capacity.
We prove, via Theorem 3.4 and A.5, that the number of parameters of any single neural pathway loaded in active GPU memory only requires $\mathcal{O}(n/\varepsilon)$ depth and $\mathcal{O}(n)$ width when approximating a Lipschitz map over the cube $[0,1]^{n}$ uniformly to an accuracy of $\varepsilon>0$. This contrasts with the optimal approximation rates for standard (non-distributed) MLP architectures, which require $\mathcal{O}(\varepsilon^{-n})$ depth and $\mathcal{O}(n)$ width to achieve the same $\varepsilon$ accuracy.
Importantly, unlike results that break the curse of dimensionality using standard MLPs, our results always hold in practice since we do not require the target function to have high regularity. Furthermore, in contrast to approximation results leveraging super-expressive activation functions with infinite VC-dimension (Proposition 3.7), any neural pathway of finite depth and width, defined by a finite tree with weight and bias constraints, has finite VC-dimension (Proposition 3.6). Thus, our models can learn and generalize effectively when trained on finite data.
We validate our algorithmic framework and mathematical analysis by testing neural pathways on regression tasks. Specifically, we focus on learning 1D Hölder functions of low regularity, as well as on 2D and 3D Ackley and Rastrigin functions. Additionally, we assess its performance in classification, particularly in higher-dimensional spaces such as the latent space of pretrained self-supervised large computer vision models like DINOv2. This not only helps verify the applicability of our model beyond synthetic experiments but also demonstrates its effectiveness in setups that are of closer interest to the applied machine learning community. In all cases, our model either matches or outperforms the centralized baseline while loading substantially lighter networks into GPU memory one at a time.
Future Work. In this paper, we present the initial implementation of our framework and we anticipate that future efforts will be necessary to scale it to larger networks. In such cases, the inclusion of deeper tree structures, akin to those discussed in our proofs, will be pertinent in addressing computational complexity. Similarly, as discussed in Section 2, our framework is modular, and we foresee that the methodology could be potentially modified depending on the downstream task and specific network architectures, as we primarily focused on MLPs only.
7 Acknowledgment and Funding
The authors would like to thank James Lucas and Rafid Mahmood for their helpful feedback while preparing the manuscript.
AK acknowledges financial support from the NSERC Discovery Grant No. RGPIN-2023-04482. HSOB acknowledges financial support from an EV travel grant.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Overview of Technical Derivation

### v2

@@ 3 Preliminaries
We use the following notation: for any $f,g:\mathbb{R}\to\mathbb{R}$, we write $f\in\mathcal{O}(g)$ if there exist $x_{0}\in\mathbb{R}$ and $M\geq 0$ such that for each $x\geq x_{0}$ we have $|f(x)|\leq Mg(x_{0})$. Similarly, we write $f\in\Omega(g)$ to denote the relation $g\in O(f)$. The ReLU activation function is given for every $x\in\mathbb{R}$ by $\text{ReLU}(x)=\max\{x,0\}$. For each $n\in\mathbb{N}_{+}$ and $C\subseteq\mathbb{R}^{n}$, the indicator function $I_{C}$ of $C$ is defined by: for each $x\in\mathbb{R}^{n}$ set $I_{C}(x)=1$ if $x\in C$ and is 00 otherwise.
3.2 Background
Multi-Layer Perceptrons
@@ 3.2 Background
We will consider MLPs with trainable PReLU activation functions.
@@
$\theta$ $\leftrightarrow\big{(}A^{(j)},b^{(j)},\bar{\alpha}^{(j)}\big{)}_{j=0}^{J}\,\,\mbox{ and }\,\,(A^{(j)},b^{(j)},\bar{\alpha}^{(j)})$ $\in\mathbb{R}^{d_{j+1}\times d_{j}}\times\mathbb{R}^{d_{j}}\times\mathbb{R}^{d_{j}}.$ (1)
We recursively define the representation function of a $[d]$-dimensional network by
@@
$x^{(j+1)}$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\sigma_{\bar{\alpha}^{(j)}}\bullet(A^{(j)}\,x^{(j)}+b^{(j)})\qquad\mbox{ for }j=0,\dots,J-1$
@@
shatters it, i.e. if all possible classifiers on $\{x_{i}\}_{i=1}^{k}$ are implementable in the sense that $\{I_{(0,\infty)}\circ f:\,f\in\mathcal{F}\}=\{0,1\}^{\{x_{i}\}_{i=1}^{k}}$; here $I_{(0,\infty)}(t)=1$ if $t>0$ and equals to 00 otherwise. Denoted by $\operatorname{VC}(\mathcal{F})$, the VC dimension of $\mathcal{F}$ is the cardinality of the largest $k$-point subset shattered by $\mathcal{F}$. If $k$ is unbounded, then we say that $\mathcal{F}$ has an infinite VC dimension (over $\mathcal{X}$). One can show, see Bartlett et al. (2019), that the VC-dimension of any such $\mathcal{F}$ is roughly the same as the pseudo-dimension of Pollard (1990) for a small modification of $\mathcal{F}$.
@@
We now define our toy mixture of experts model, the MoMLP.
3.3 Definition: Our Toy Mixture of Experts Model - the MoMLP
Figure 1: $1$) The distance from each input $x$ to all prototypes $p_{1},\dots,p_{8}$ ($\ell=8$) is queried. $2$) The network ($\hat{f}_{2}$ in the figure) assigned to the nearest prototype ($p_{2}$), is loaded onto the GPU and used for prediction.
We study the following toy MoE model, where each expert (P)ReLU MLP specializes in a distinct region of the input domain $[0,1]^{n}$. Informally, these regions $C_{1},\dots,C_{\ell}$ correspond to the sets of closest points (Voronoi cells) from a finite set of prototypes/landmarks $p_{1},\dots,p_{\ell}$ in $[0,1]^{n}$, as illustrated in Figure 1. Associated to each region $C_{i}$ is a single expert MLP $\hat{f}_{i}$ with (P)ReLU activation function responsible for approximating the target function only thereon. Here, the sparse gating procedure which routes any given input $x\in[0,1]^{n}$ to the expert corresponding to the nearest prototype $p_{i}$ is implemented by a (finite) routing tree $\mathcal{T}=(V,E)$ whose nodes $V$ are points in $[0,1]^{n}$ and leaves (terminal nodes) are the points $p_{1},\dots,p_{\ell}$.
We now formally define the classes of MoMLPs.
Definition 3.2 (MoMLPs).
Let $J,W,{l},n\in\mathbb{N}$ and fix an activation function $\sigma\in C(\mathbb{R})$. The set of MoMLPs with at-most $L$ leaves, depth $J$, and width $W$, denoted by $\mathcal{NP}_{J,W,{l}:n,m}^{\sigma}$, consists of all functions $\hat{f}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ satisfying $\hat{f}=\sum_{i=1}^{l}\,f_{i}\,I_{C_{i}}$ where for $f_{1},\dots,f_{l}\in\mathcal{NN}^{\operatorname{PReLU}}_{J,W}$, and distinct prototypes $p_{1},\dots,p_{{l}}\in\mathbb{R}^{n}$; whose induced Voronoi cells $\{C_{i}\}_{i=1}^{{l}}$ are defined recursively by
$C_{i}$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\tilde{C}_{i}\setminus\bigcup_{j<i}\,\tilde{C_{j}},$ (5)
where for $i=1,\dots,l$ the (non-disjoint) cells are
$\tilde{C}_{i}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\big{\{}x\in[0,1]^{n}:\,\|x-p_{i}\|=\min_{j\in\{1,\dots,l\}}\,\|x-p_{j}\|\big{\}}.$ (6)
The Routing Trees
The structure in any MoMLP is any tree with root node $\mathbb{R}^{n}$ and leaves given by the pairs $\{(p_{i},f_{i})\}_{i=1}^{l}$ or equivalently $\{(C_{i},f_{i})\}_{i=1}^{l}$. The purpose of any such tree is simply to efficiently route an input $x\in\mathbb{R}^{n}$ to one of the $l$ “leaf networks” (the experts) $f_{1},\dots,f_{l}$ using $\mathcal{O}(l\log(l))$ queries; to identify which Voronoi cells $\{C_{i}\}_{i=1}^{l}$ the point $x$ is contained in.
Remark 3.3 (Partitioning in (5) in classical computer science).
4 Main Results
We first present our main approximation theoretic guarantee, which gives complexity estimates for mixtures of MLPs with trainable PReLU activation functions when uniformly approximating arbitrary locally-Hölder function on the closed unit ball of $\mathbb{R}^{n}$, defined by $\overline{B}_{n}(0,1)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{x\in\mathbb{R}^{n}:\,\|x\|\leq 1\}$.
Our rates depend on a “number of experts-to-expert complexity trade-off parameter” $r\in\mathbb{R}$ which determines how fast the overall MoE complexity scales, in terms of the number of experts and the complexity of each expert, as the approximation error becomes small. Setting $r<0$ implies that more model complexity will be loaded into each expert MLP and there will be fewer experts defining the MoE. In contrast, setting $r>0$ loads less complexity in each expert MLP at the cost of more experts in the MoE. In particular, when $r=0$, each expert will have constant complexity even when the approximation error becomes arbitrarily small.
Theorem 4.1 (Trade-Off: No. Expert vs. Expert Complexity).
Suppose that $\sigma$ satisfies Definition 3.1. Fix an “number of experts-to-expert complexity trade-off parameter” $r\in\mathbb{R}$.
@@ 4 Main Results
For every $\alpha$-Hölder map $f:\overline{B}_{n}(0,1)\to\mathbb{R}^{m}$ with $0<\alpha\leq 1$ and each approximation error $\varepsilon>0$, there is a $p\in\mathbb{N}_{+}$, a binary tree $\mathcal{T}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(V,E)$ with leaves =$\mathcal{L}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{(v_{i},\theta_{i})\}_{i=1}^{L}\subseteq{\mathcal{K}}\times\mathbb{R}^{p}$ and a family of MLPs with (P)ReLU activation function $\{\hat{f}_{\theta_{i}}\}_{i=1}^{L}$ defined by $p$ parameters and mapping $\mathbb{R}^{n}$ to $\mathbb{R}^{m}$ satisfying:
$\max_{x\in{\mathcal{K}}}\,\min_{(v_{i},\theta_{i})\in\mathcal{L}}\,\|x-v_{i}\|\in\Theta\big{(}\varepsilon^{1/\alpha-2r/n}\big{)}$
and for each $x\in{\mathcal{K}}$ and $i=1,\dots,L$, if $\|x-v_{i}\|<\delta$ then
$\|f(x)-f_{\theta_{i}}(x)\|<\varepsilon.$
@@
Next, we demonstrate that the MoMLP model can generalize and generate functions that are PAC-learnable, thanks to its finite VC dimension. This property, however, breaks down in MLP models with super-expressive activation functions.
Theorem 4.2 (VC-Dimension Bounds for MoMLPs - MoMLPs Can Generalize).
@@
4.1 Discussion
Trade-off between Number of Experts and Expert Complexity. Our results suggest that, theoretically, successful MoE models may not need each expert to be highly overparameterized if there are enough experts. This hypothesis is ablated experimentally in Appendix D in the context of irregular function approximation in low-dimension space; which is equivalent to high-dimensional regular function approximation (see Appendix C for a discussion on this later point).
Logarithmic number of queries via trees. For many prototypes, as in our main guarantee, the MoMLPs only need to evaluate the distance between the given input and a logarithmic number of prototypes—specifically, one for each level in the tree—when using deep binary trees to hierarchically refine the Voronoi cells. Thus, a given machine never processes the exponential number of prototypes, and only $\nu\lceil\log_{\nu}(K)\rceil$ prototypes are ever queried for any given input; when trees are $\nu$-ary (as in Theorem 5.3), and where $K$ denotes the number of prototypes. Since we consider that prototypes are queried separately and before loading MoMLPs, we do not take them into account when counting the number of learnable parameters. Moreover, the size of our prototypes is negligible in our experiments.
4.2 Application: Controlling The Complexity in VRAM maintaining a Finite VC Dimension
Super-Expressive Activation Functions Have Infinite VC-Dimension. We complement the main result of Bartlett et al. (2019) by demonstrating that the class of unstable MLPs (Shen et al., 2022a) possesses infinite VC dimension, even when they have finite depth and width. Thus, while they may serve as a gold standard from the perspective of approximation theory, they should not be considered a benchmark gold standard from the viewpoint of learning theory.
Definition 4.3 (Trainable Super-Expressive Activation Function).
where $\sigma^{\star}:\mathbb{R}\to\mathbb{R}$ is given by
@@ 4.2 Application: Controlling The Complexity in VRAM maintaining a Finite VC Dimension
Let $\mathcal{F}$ denote the set of MLPs with activation function in Definition 4.3, depth at-most $15$, and width at-most $36n(2n+2)$. Then $\operatorname{VC}(\mathcal{F})=\infty$.
@@
The VC dimension bounds for the standard MLP model, MLP with a super-expressive activation function as proposed by Shen et al. (2022a), and the MoMLP model are summarized in Table 2.
@@
Table 2: VC Dimension of the MoMLPs, ReLU MLP, and MLP model with Super-Expressive Activation function of Shen et al. (2022a). All models have depth $J$, width $W$, and (when applicable) $L$ leaves; where $J,W,L,n\in\mathbb{N}_{+}$.
Model VC Dim Ref.
MoMLPs $\mathcal{O}\big{(}L\log(L)^{2}\,\max\{nL\log(L),JW^{2}\log(JW)\}\big{)}$ Thrm 4.2
ReLU MLP $\mathcal{O}\big{(}JW^{2}\log(JW)\big{)}$ Bartlett et al. (2019)
Super-Expressive $\infty$ Prop 4.4
5 Overview of Derivation